OpenStudy (anonymous):

simplify this 7a-4a(b+3)

OpenStudy (anonymous):

=7a-4ab-12a =-5-4ab =-4ab-5

OpenStudy (anonymous):

7a - 4ab - 12a -5a - 4ab